a fixed quantity of gas is compressed at constant tempature from a volume of 638 mL to 208 mL. if the initial pressure was 522 mm Hg, what is the final pressure?
1
Expert's answer
2020-06-02T14:04:05-0400

Solution: According to law Mendeleev-Clapeyron at constant temperature: P1V1=P2V2. We known that P1=522 mm Hg and V1=638 mL are inital units, and V2=208 mL is a final volume, then P2=P1V1/V2=522*638/208=1601 mm Hg is the final pressure


Answer: 1601 mm Hg.